The length of the first side of a triangle is twice the length of the second side, and the length of the third side is 5 more than the length of the second side. Express the perimeter of the triangle in terms of the length of its second side, and use this model to estimate the lengths of the three sides if the perimeter is to be 100 in.
Answer by mananth(16946)

The length of the first side of a triangle is twice the length of the second side, and the length of the third side is 5 more than the length of the second side. Express the perimeter of the triangle in terms of the length of its second side, and use this model to estimate the lengths of the three sides if the perimeter is to be 100 in.
...
second side = x
first side= 2x
third side= x+5
...
x+2x+x+5======> 4x+5 = Perimeter
...
4x+5 =100
4x=95
/4
x=23.75 in
2x= 47.50 in
x+5 = 28.75 in